# Monitored long-range interacting systems: spin-wave theory for quantum trajectories

### v1

@@ III Spin-wave theory along quantum trajectories
In this section, we formulate the theory of non-equilibrium spin-wave quantum trajectories (SWQT) and derive the equations of motion for spin systems under continuous monitoring. This formalism serves as a semi-classical method for solving quantum trajectories of generic out-of-equilibrium dissipative spin systems with sufficiently long-range interactions whose average dynamics can be described by a Lindblad master equation. In previous studies [97], the non-equilibrium spin-wave theory has been generalized to dissipative spin systems on the level of the averaged state, i.e. the density matrix evolving deterministically under the master equation, cf. Eq. (6). This technique has been proven successful in studying dissipative dynamics of large $N$ over extended time spans. Our method, on the other hand, resolves the dynamics on the level of single quantum trajectories, corresponding to the unravelling (1) of the master equation, enabling us to probe non-linear quantum-trajectory correlations, encoded in the entanglement and other non-linear functions of the state [107].
The central assumption is that the system admits a strong collective spin polarization, on top of which spin-wave excitations are bosonized via a truncated Holstein-Primakoff expansion. The bosonic modes are then approximated with a Gaussian ansatz [108, 109] parametrized by first and second moments. We remark that, since our assumption is formulated at the level of single trajectories and not at the level of average dynamics, our variational ansatz allows us to reach a larger class of density matrices compared to previous methods [109]. Indeed, the mixture of Gaussian states, such as those describing single trajectories, is a non-Gaussian state in general, hence encodes non-trivial correlations between the spin-waves. This fact has a fundamental operational consequence: spin-wave quantum trajectories provide more accurate representations of the purely Lindblad dynamics [109], making the SWQT compelling also for the study of the dissipative dynamics of hermitian operators.
The main challenge in the generalization of spin-wave theory to quantum trajectories comes from the fact that the state no longer follows a smooth deterministic evolution, as the density matrix does, but rather evolves stochastically, whereas the bosonization of spin-wave excitations has to be performed in the instantaneous frame aligned with the collective spin polarization in order to yield a valid approximation. In the case of interest, the evolution is continuous but nowhere differentiable. Therefore, a naive rotating frame cannot be directly applied. Therefore, we base our approach on the Euler–Maruyama [110] method to determine the coupled stochastic evolution of the state and of the frame in a self-consistent way.
The rest of the section is organized as follows. In Sec. III.1, we introduce the major approximations made in the spin-wave theory, namely the truncated Holstein-Primakoff expansion and the Gaussian approximation. Then, We derive the equations of motion in Sec. III.2.
III.1 Spin-wave approximation
To keep the discussion self-consistent, we introduce in this subsection the key approximations made in our spin-wave theory, namely the truncated Holstein-Primakoff transformation and the Gaussian approximation. These approximations will bosonize the spin fluctuations on top of a moving vacuum defined by the collective polarization of all spins, and then approximate the bosonic fluctuations up to the Gaussian, i.e. quadratic, level.
Figure 1: (a) Sketch of the reference frames considered in our method. $Oxyz$ is the lab frame and $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ represents the instantaneous frame where ${\tilde{z}}$ aligns with the collective mode polarization. For illustration purposes, the axes ${\tilde{x}}$ and ${\tilde{y}}$ are moved from the origin to the surface of the sphere for clarity. (b) Sketch of the re-alignment of the reference frame where the frame angles $(\theta,\phi)$ are updated to match the collective spin direction after its evolution in an infinitesimal time step. The curved arrows represent the paths $\gamma_{1}$ and $\gamma_{2}$ as defined in Eq. (36) for evaluating the integral of Eq. (35). The dashed (solid) straight arrow represents the ${\tilde{z}}$ axis before (after) the re-alignment step.
To prepare for this discussion, let us now consider a generic rotated reference frame $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ defined by the unitary transformation parametrized by angles $\theta$ and $\phi$ [see Fig. 1 (a) for an illustration of the rotated frame]
$\hat{U}(\theta,\phi)={\rm{e}}^{-{\rm{i}}\phi\hat{S}^{z}}{\rm{e}}^{-{\rm{i}}\theta\hat{S}^{y}}\,.$ (16)
The spin operators associated with the rotated frame are given by
$\hat{s}_{i}^{\tilde{\alpha}}=\hat{U}(\theta,\phi)\hat{s}_{i}^{\alpha}\hat{U}^{\dagger}(\theta,\phi)=\sum_{\beta}G_{{\tilde{\alpha}}\beta}\hat{s}^{\beta}_{i}\,,$ (17)
where $G_{{\tilde{\alpha}}\beta}\equiv\vec{e}_{{\tilde{\alpha}}}\cdot\vec{e}_{\beta}$ is the $3\times 3$ matrix representing the spatial rotation
$G=\begin{pmatrix}\cos\theta\cos\phi&\cos\theta\sin\phi&-\sin\theta\\ -\sin\phi&\cos\phi&0\\ \sin\theta\cos\phi&\sin\theta\sin\phi&\cos\theta\end{pmatrix}\,.$ (18)
Here, the rows of $G$ are the rotated unit vectors $\bm{e}_{\tilde{x}}$, $\bm{e}_{\tilde{y}}$ and $\bm{e}_{\tilde{z}}$ (assuming $\{\bm{e}_{\alpha}\}_{\alpha}$ to be the canonical basis for the lab frame). The inverse transformation
$\hat{s}_{i}^{\beta}=\sum_{{\tilde{\alpha}}}G_{{\tilde{\alpha}}\beta}\hat{s}^{{\tilde{\alpha}}}_{i}\,,$ (19)
allows us to express all the operators in the rotated frame. We now choose the frame angles $(\theta,\phi)$ such that the rotated ${\tilde{z}}$ aligns with the direction of the collective spin
$\langle\hat{S}^{\tilde{x}}\rangle=\langle\hat{S}^{\tilde{y}}\rangle=0\,.$ (20)
This allows us to approximate the spin operators with a Holstein-Primakoff transformation truncated to the lowest order
$\hat{s}^{{\tilde{z}}}_{i}$ $=s-\hat{b}^{\dagger}_{i}\hat{b}_{i}\,,$ (21)
$\hat{s}^{{\tilde{x}}}_{i}$ $\simeq\sqrt{\dfrac{s}{2}}(\hat{b}^{\dagger}_{i}+\hat{b}_{i})\,,$
$\hat{s}^{{\tilde{y}}}_{i}$ $\simeq{\rm{i}}\sqrt{\dfrac{s}{2}}(\hat{b}^{\dagger}_{i}-\hat{b}_{i})\,,$
@@ III.1 Spin-wave approximation
where $\hat{b}_{i}$’s are bosonic operators with standard bosonic commutation relations $[\hat{b}_{i},\hat{b}^{\dagger}_{j}]=\delta_{ij}$. This transformation effectively approximates the Bloch sphere with the tangent plane at the north pole of the rotated frame [2], which is why we need to align ${\tilde{z}}$ with the collective spin direction. This approximation becomes exact only when $\langle\hat{b}^{\dagger}_{i}\hat{b}_{i}\rangle=0$, i.e., when the system is in a spin coherent state. Thus, the density of bosonic excitations
@@
serves as a natural control parameter for the approximation. In the regime of $\epsilon\ll 1$, the bosonization in Eq. (21) faithfully captures the spin dynamics without higher-order corrections [97]. The bosonic modes $\hat{b}_{i}$ represent spin-wave excitations on top of the collective polarization. We will therefore refer to the control parameter $\epsilon$ as the spin-wave density.
Let us now make our second approximation, where we assume the state of the bosonic modes to be Gaussian along each trajectory. This allows us to uniquely specify the state of the entire system with the following variational parameters:
@@
where $\hat{\delta}_{i}\equiv\hat{b}_{i}-\beta_{i}$. Therefore, this method requires only $\mathcal{O}(N^{2})$ parameters to represent the state in the most general case, which is exponentially efficient as compared to the dimension of the full Hilbert space $2^{N}$. The complexity can be further reduced if additional symmetry is present in the dynamics. For example, the power-law spin model introduced in Sec. II.1 preserves translational invariance on the level of single trajectories. Therefore, the covariances $u_{ij}$ and $v_{ij}$ will depend only on the distance between sites $i$ and $j$, which reduces the number of parameters to $\mathcal{O}(N)$.
The Gaussian approximation, as a variational ansatz, allows us to calculate any physical quantity of the state within the approximation. In particular, the entanglement between spins can be readily obtained using the covariance formalism of Gaussian states [111], see Appendix C for details.
III.2 Equations of motion
In order to derive the unraveled dynamics under the spin-wave approximation introduced in the previous subsection, we should provide the rules for updating the variational parameters of the state, $\theta$, $\phi$, $\beta_{i}$, $u_{ij}$ and $v_{ij}$ according to the stochastic unraveling defined in Eq. (1). Our algorithm is based on the Euler–Maruyama method [110], where time is discretized into small steps. Initializing the parameters at $t=0$ such that the rotated frame parametrized by $(\theta,\phi)$ has its ${\tilde{z}}$ axis aligned with the collective spin (which implies $\sum_{i}\beta_{i}=0$), we proceed stroboscopically by repeating the two following steps until the desired time $t$ is reached:
@@ III.2 Equations of motion
    Update frame angles $(\theta,\phi)$ such that $\sum_{i}\beta_{i}=0$ in the new rotated frame [see Fig. 1 (b) for a schematic representation of this step]. This can be achieved self-consistently by considering the evolution of the bosonic mode $\hat{b}_{i}$ generated by the (passive) rotation of the frame alone. Then update the Gaussian parameters $\beta_{i}$, $u_{ij}$ and $v_{ij}$ accordingly (due to the rotation of the frame). Finally, increase the time by $\delta t$ and start a new iteration.
The re-alignment of the reference frame $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ in step 2 above is essential to the validity of the truncated Holstein-Primakoff bosonization approximation (which is only sound for small deviations of the spin from the north pole of the Bloch sphere). This will allow the accurate representation of the dynamics over an extended time window significantly beyond the restrictions of a conventional spin-wave theory bosonizing in a static reference frame.
We detail below the key ingredients for performing the two steps sketched above.
III.2.1 Infinitesimal increments

### v2

@@ II.2 Spin-wave theory along quantum trajectories
In this section, we present the framework of spin-wave quantum trajectories (SWQT), that is the main result of our work. This formalism serves as a semi-classical method for solving quantum trajectories of generic out-of-equilibrium dissipative/monitored spin systems with sufficiently long-range interactions whose average dynamics can be described by a Lindblad master equation. The resolution of single quantum trajectories enables us to probe non-linear correlations, encoded in the entanglement and other non-linear functions of the state [64].
The central assumption is that the system admits a strong collective spin polarization, on top of which spin-wave excitations are bosonized via a Holstein-Primakoff expansion truncated to the lowest order:
$\hat{s}^{{\tilde{z}}}_{i}(\theta,\phi)$ $=s-\hat{b}^{\dagger}_{i}\hat{b}_{i}\,,$ (9)
$\hat{s}^{{\tilde{x}}}_{i}(\theta,\phi)$ $\simeq\sqrt{\dfrac{s}{2}}(\hat{b}^{\dagger}_{i}+\hat{b}_{i})\,,$
$\hat{s}^{{\tilde{y}}}_{i}(\theta,\phi)$ $\simeq{\rm{i}}\sqrt{\dfrac{s}{2}}(\hat{b}^{\dagger}_{i}-\hat{b}_{i})\,.$
@@
Here, the $\hat{s}^{{\tilde{\alpha}}}$’s are spin operators in the rotated frame $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ as illustrated in Fig. 1 (b). They are related to the lab-frame spin operators via $\hat{s}_{i}^{\tilde{\alpha}}(\theta,\phi)=\hat{U}(\theta,\phi)\hat{s}_{i}^{\alpha}\hat{U}^{\dagger}(\theta,\phi)$ with $\hat{U}(\theta,\phi)={\rm{e}}^{-{\rm{i}}\phi\hat{S}^{z}}{\rm{e}}^{-{\rm{i}}\theta\hat{S}^{y}}$. The ${\tilde{z}}$ axis of the rotated frame is aligned with the collective (average) polarization of all the spins, which can be met by imposing $\langle\hat{S}^{\tilde{x}}\rangle=\langle\hat{S}^{\tilde{y}}\rangle=0\,$. The $\hat{b}_{i}$’s are bosonic operators, representing spin fluctuations around the collective magnetization, with standard bosonic commutation relations $[\hat{b}_{i},\hat{b}^{\dagger}_{j}]=\delta_{ij}$. This transformation effectively approximates the Bloch sphere with the tangent plane at the north pole of the rotated frame [2], and becomes exact only when $\langle\hat{b}^{\dagger}_{i}\hat{b}_{i}\rangle=0$, i.e., when the system is in a spin coherent state. Thus, the density of bosonic excitations
@@
also referred to as spin-wave density, serves as a natural control parameter for the approximation [58].
The bosonic modes are then approximated with a Gaussian ansatz [65, 66] parametrized by first and second moments. This allows us to uniquely specify the state of the entire system with the following variational parameters:
@@
where $\hat{\delta}_{i}\equiv\hat{b}_{i}-\beta_{i}$. Therefore, this method requires only $\mathcal{O}(N^{2})$ parameters to represent the state in the most general case, which is exponentially efficient as compared to the dimension of the full Hilbert space $2^{N}$, and the complexity can be further reduced in the presence of additional symmetries.
The stochastic master equation (1) then translates into the dynamics of the variational parameters (11). Our algorithm is based on the Euler–Maruyama method [67], where time is discretized into small steps $\delta t$. Initializing the parameters at $t=0$ such that the rotated frame parametrized by $(\theta,\phi)$ has its ${\tilde{z}}$ axis aligned with the collective spin (which implies $\sum_{i}\beta_{i}=0$), we proceed stroboscopically by repeating the two following steps until the desired time $t$ is reached:
@@
    Update frame angles $(\theta,\phi)$ such that $\sum_{i}\beta_{i}=0$ in the new rotated frame [see Fig. 1 (c) for a schematic representation of this step]. This can be achieved self-consistently by considering the evolution of the bosonic mode $\hat{b}_{i}$ generated by the (passive) rotation of the frame alone. Then update the Gaussian parameters $\beta_{i}$, $u_{ij}$ and $v_{ij}$ accordingly (due to the rotation of the frame). Finally, increase the time by $\delta t$ and start a new iteration.
We detail in Sec. IV.1 the key ingredients for performing the two steps sketched above. Note that as this construction does not depend on the spatial structure of the system, it is straightforward to apply the method to arbitrary dimensions and to adapt to different lattice geometries, without comprising its computational efficiency. Moreover, we would like to stress that the framework constructed here is fully generic, as it suffices to systematically apply the bosonization and the Gaussian approximation to obtain the equations of motion. Therefore, it can be straightforwardly extended to other types of systems beyond those we considered. An example of its generalization to spin-boson systems is presented in Supplementary Note E together with some illustrative numerical results.
Finally, let us remark that since our approximations are formulated at the level of single trajectories and not the average dynamics, our variational ansatz allows us to reach a larger class of density matrices [66] compared to previous methods working on the level of the density matrix [58]. Indeed, the mixture of Gaussian states, such as those describing single trajectories, is a non-Gaussian state in general, hence encodes non-trivial correlations between the spin-waves. This fact has a fundamental operational consequence: spin-wave quantum trajectories provide more accurate representations of the purely Lindblad dynamics, making the SWQT framework compelling also for the study of the dissipative dynamics of hermitian operators.
II.3 Entanglement phase transition in the power-law spin model
In this section, we apply the presented framework of spin-wave quantum trajectories to study the power-law interacting spin model introduced in Sec. II.1.1. We focus on $s=1/2$ and fix the interaction strength at $J=0.1\kappa$ for the rest of the section.
II.3.1 Infinite-range case: $\alpha=0$
We now test the validity of the approximations by evaluating both linear (e.g., observables) and non-linear (e.g., entanglement entropy) functions of the state. In order to benchmark the method, we first revisit the known case of $\alpha=0$, where the $z-z$ interaction is all-to-all with no spatial resolution. In this regime, the system can be effectively represented as a single spin $S=N/2$, allowing us to benchmark the spin-wave method against the exact simulation of the dynamics in the Dicke basis.
Figure 2: Benchmark of the spin-wave method against the exact solution for the time-evolution of the following quantities: (a) the magnetization $\langle\hat{S}^{z}\rangle$ along a single trajectory; (b) the half-system entanglement entropy $S_{E}$ along the same trajectory as in (a); (c) trajectory-average of the collective spin vector $\overline{\langle\hat{S}^{\alpha}\rangle}$; (d) trajectory-average of the half-system entanglement entropy $\overline{S_{E}}$. In single-trajectory benchmarks [(a) and (b)], the same noise realization is adopted by both the spin-wave quantum trajectory and the exact integration of the stochastic master equation. The trajectory-average benchmarks [(c) and (d)] are performed over 4000 trajectories. Parameters: $\omega=1.25\kappa$, $J=0.1\kappa$, $S=64$, $\kappa\delta t=10^{-4}$.
To put the spin-wave quantum trajectories in fair comparison with the exact ones, we integrate the exact stochastic master equation (1) using the Euler-Maruyama method, adopting the same time step and the same noise realization as used in the spin-wave calculation. An example of a single-trajectory benchmark for the spin-wave method is shown in Fig. 2 (a) and (b), where we compare trajectories for both the magnetization $\langle\hat{S}^{z}\rangle$ and the half-system entanglement entropy (cf. Supplementary Note C and Ref. [68, 69]) $S_{E}$ against exact ones. The considered system has $S=64$ and $\omega=1.25\kappa$, which, as predicted by the mean-field theory, corresponds to the time-crystal phase in the thermodynamical limit. The initial state is the fully polarized Dicke state (all spins pointing up) for both simulations. Surprisingly, the trajectories obtained with the spin-wave method faithfully reproduce the dynamics of both quantities throughout most of the evolution. This suggests that the proposed method can be used as an approximation to resolve dynamics on the level of single trajectories.
We also evaluate the performance of the spin-wave method on trajectory-averaged quantities. Fig. 2 (c) shows the time-evolution of the spin expectations $\overline{\langle\hat{S}^{x,y,z}\rangle}$ obtained by averaging spin-wave quantum trajectories. These are then benchmarked against the exact solution of the Lindblad master equation (8). As expected from the single-trajectory performance, the average dynamics is accurately reproduced. Notably, the decaying oscillations of the magnetization are due to the finite size effect and therefore require corrections beyond the mean-field level to capture. It is interesting to note that the length of the averaged collective spin vector $\overline{\langle\hat{S}^{\alpha}\rangle}$ is only a fraction of the maximal value $S$, which implies the that the average density matrix, of highly mixed nature, is far from a state that can be well approximated by the spin-wave approximations. As a consequence, a deterministic version of the spin-wave theory with the same approximations applied to the averaged state, e.g., the one proposed in Ref. [58], will completely fail to capture the dynamics in this regime, cf. Supplementary Note D.2. In Fig. 2 (d), we compare the trajectory-averaged entanglement entropy with that calculated from exact trajectories, which shows a good agreement again. We also report the average spin-wave density to be $\overline{\epsilon}\lesssim 10^{-2}$ throughout the time evolution (see Supplementary Note D.1 for a detailed discussion and additional benchmarks).
Figure 3: Results for the power-law interacting model in the long-range regime ($\alpha=0.2$, $J=0.1\kappa$) of different system sizes (colorbar shared across panels). The dashed line marks the mean-field solution. (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. (b) Steady-state of the trajectory-averaged half-chain entanglement entropy as a function of $\omega$. The vertical dotted line marks the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. Inset: scaling of the maximum entropy versus $N$ in linear-log scale.
Figure 4: Maximum steady-state entanglement as a function of system size $N$ (total number of spins) for different interaction range $\alpha$ in (a) 1D and (b) 2D. Different scaling behavior can be observed (see annotation). Note that the horizontal axes are truncated for better visualization and the maximum number of spins we report is $N=1024$ in 2D (see full figures in Supplementary Note D.5). We report that the value of the drive $\omega$ maximizing the entanglement is asymptotically independent of $N$ (cf. Supplementary Note D.5 Fig. S8.)
II.3.2 Long-range case: $\alpha\neq 0$
To investigate the effect of a finitely long-range interaction and the entanglement dynamics in this regime, we first apply the spin-wave method to the model in one dimension (1D) with $\alpha=0.2$. Fig. 3 (a) shows the steady-state expectation of the collective magnetization $\overline{\langle\hat{S}^{z}\rangle}$ as a function of the drive amplitude $\omega$ and the system size $N$. As we approach the thermodynamic limit by increasing $N$, the spin-wave solution converges towards the mean-field prediction and a continuous transition emerges separating the normal phase with $\overline{\langle\hat{S}^{z}\rangle}\neq 0$ and the time-crystal phase with $\overline{\langle\hat{S}^{z}\rangle}=0$. Fig. 3 (b) shows the behavior of the long-time-averaged half-chain entanglement entropy across this dissipative phase transition. The entanglement develops a logarithmic divergence as a function of $N$ at the critical point (as implied by the finite-size scaling analysis of the maximum value of the entanglement entropy), while it appears to be $N$-independent for drive values deep inside both phases. This suggests the emergence of an entanglement phase transition separating two area-law phases. We also report (cf. Supplementary Note D.3) that the steady-state spin-wave density $\overline{\epsilon}$ is suppressed with increasing $N$, implying the asymptotic exactness of our results. Due to the sufficiently long-range interaction ($\alpha=0.2$), these results bear similar features to the infinite-range model ($\alpha=0$). For larger values of $\alpha$, the entanglement dynamics can be dramatically modified, while the dissipative transition witnessed by the magnetization remains however qualitatively similar since the transition is driven by the infinite-range dissipation and not the long-range coherent interaction. We focus therefore on the entanglement in the discussion that follows. As shown in Fig. 4 (a), the maximum entanglement for $\alpha=1.0$ in 1D exhibits a volume-law scaling in contrast to the log scaling for $\alpha=0.2$ discussed above. (See Supplementary Note D.4 for additional results in this regime.)
We also perform simulations in 2D (for up to $32\times 32=1024$ long-range interacting spins) to study the entanglement scaling as a function of $\alpha$, as shown in Fig. 4 (b) for a square lattice of size $N=L^{2}$. As we shorten the interaction range by increasing $\alpha$, the entanglement scaling changes from the log law to area law ($\propto L$) and volume law ($\propto L^{2}$), which suggests the presence of a phase transition driven by $\alpha$ between the different entanglement scaling behavior. We report (cf. Supplementary Note D.5) that even in the volume-law phase under moderately short-range settings ($\alpha=2.0$ and $\alpha=3.0$), the spin-wave density remains small at $\overline{\epsilon}\lesssim 0.1$. On the other hand, the increase of the spin-wave density with higher values of $\alpha$ suggests the eventual breakdown of the spin-wave approximations in the presence of sufficiently short-range interactions.
II.4 Experimental observability of monitored phases boosted by spin-wave quantum trajectories
In view of the experimental implementations, for example on digital quantum simulators with discrete-time dynamics, we consider a discretized version of the monitored dynamics, where the noise is binned and approximated with a binary random variable, realized via a set of ancilla qubits in the weak measurement formalism (see Sec. IV.2 for a detailed derivation). We consider an experiment with $\mathcal{M}\gg 1$ weak measurement steps (i.e. projective measurements on the ancilla) and set the measurement outcome history ${\bm{m}}$. Suppose we perform a final projective measurement of a system operator $\hat{O}$ over $\hat{\rho}({{\bm{m}}})$, i.e. the state conditioned on the measurement history ${\bm{m}}$. This measurement is disruptive and will collapse $\hat{\rho}$ onto the eigenspace of the measurement outcome $o_{{\bm{m}}}$. For example, in the case where the observable is chosen to be $\hat{O}=\hat{S}^{z}$, the final single-shot measurement outcome $o_{\bm{m}}\in\{-S,\cdots,S\}$ is one of the $12S+$ 1 eigenvalues of operator $\hat{S}^{z}$. Averaging over this quantity recasts the Lindblad prediction, namely $\overline{o_{{\bm{m}}}}={\rm{Tr}}[\hat{\rho}\hat{O}]$ where $\hat{\rho}=\overline{\hat{\rho}({\bm{m}})}$ is the average state over all possible trajectories ${\bm{m}}$. A correlation that is non-linear in the state can be obtained using classical simulations [44]. Fixing the measurement history and position ${\bm{m}}$ in the classical simulation, we obtain the (classical) estimate of the trajectory-wise expectation value $\langle\hat{O}_{{\bm{m}}}\rangle_{C}\equiv{\rm{Tr}}[\hat{\rho}_{C}({\bm{m}})\hat{O}]$ with the label $\bullet_{C}$ denoting the classically computed quantity. For instance, this is obtained in our setup using the solution of Eq. (1) within the spin-wave approximations. We can then cross-correlate the measurement outcome with the classical simulation result to construct the quantum-classical object $o_{\bm{m}}\langle\hat{O}_{{\bm{m}}}\rangle_{C}$, which, when averaged over ${\bm{m}}$ (i.e. over measurement shots),
gives the quantum-classical correlator, where the label $\bullet_{Q}$ denotes the quantity evaluated on the actual quantum state in the experiment. In the ideal case where the classical simulation is exact, the quantum and classical quantities should coincide, giving a nonlinear function of the state conditioned on the measurement history. This was recently shown to reproduce the measurement-induced transition in a variety of experiments [31], provided simulating $\langle\hat{O}_{\bm{m}}\rangle_{C}$ is easy. The SWQT framework therefore is a compelling toolbox for investigating quantum-classical correlations. In particular, when the semi-classical approximation holds, we expect the data to be reliable with errors $O(1/S)$.
Figure 5: Steady-state trajectory-averaged quantum-classical correlator $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{C}}$ as a function of the drive $\omega$ for different total spin numbers $S$ (see legend) for the infinite-range model ($\alpha=0$ and $J=0.1\kappa$). The vertical dotted line marks the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. The steady-state value is computed as the average from $\kappa t=50$ to $\kappa t=100$, i.e. over the tail of the dynamics.
To demonstrate this quantum-classical measurement protocol, we revisit the infinite-range ($\alpha=0$) case of the power-law spin model (fixing again $J=0.1\kappa$), where the numerically exact solution, that we use to mimic the “quantum” run, is affordable. The ‘classical” quantities are then those given by the SWQT method using the same noise realization ${\bm{m}}$. In Fig. 5, we show the steady-state behavior of the quantum-classical correlator $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{C}}$ as a function of the drive $\omega$ for different system sizes. Importantly, this quantity successfully captures the signature of the phase transition. This demonstrates that the efficient classical simulation enabled by our spin-wave framework provides access to quantum-classical observables, which can probe nonlinear properties of the state without the post-selection issue.
III Discussion
In this work, we have proposed a stochastic spin-wave theory along quantum trajectories of monitored long-range spin systems. Our results illustrate several features of the framework that can be summarized as follows:
  • The methodology developed here has a wide applicability not only in the study of entanglement phase transitions but can also be exploited to calculate any other quantity, linear and non-linear in the quantum state along the dynamics either on the level of trajectories or on the Lindblad level, by taking trajectory averages.
  • The proposed method works equally well in any dimension, namely its complexity is always polynomial in the total number of qubits. This overcomes the typical limitation of tensor network methods in two or higher dimensions, including volume-law scaling phases.
  • When the semi-classical approximations are valid – including the relevant case of sufficiently long-range interactions, the efficient classical simulation of spin-wave quantum trajectories allows to probe nonlinear quantities of the state via quantum-classical cross-correlated observables, which does not suffer from the post-selection problem, thus opening up promising avenues in the experimental detection of monitored phases in long-range systems.
Despite the demonstrated performance of our method, there are several interesting directions for future improvements. For example, one can consider terms in the Holstein-Primakoff expansion beyond the leading order to include more nonlinear effects, which should extend the validity of the theory in the regime of relatively high spin-wave densities. We also expect better overall single-trajectory accuracy with the inclusion of higher-order spin-wave corrections, which is essential in improving the performance of the quantum-classical measurement protocol in avoiding the post-selection problem. On the other hand, different unravellings of the Lindblad master equation can also be explored, including the relevant case of quantum jumps modeling the coupling to photodetectors.
The long-range spin model we adopted for the demonstration of the spin-wave theory can be readily realized in current experimental platforms [71, 8], and we expect the entanglement phase transition to be readily observable in experiments without the post-selection problem.
IV Methods
IV.1 Equations of motion for spin-wave quantum trajectories
We provide in this section the technical details for performing the two steps in evolving the variational state ansatz as sketched in Sec. II.2.
IV.1.1 Infinitesimal increments